Please solve: 75=.0775x+x
Found 2 solutions by AnlytcPhil, tutorcecilia:
Answer by AnlytcPhil(1806)

Please solve:

           75 = .0775x + x

Remove the decimal by multiplying every 
term by 10,000

       750000 = 775x + 10000x

Combine the two like terms on the right

       750000 = 10775x

Divide both sides by 10775

 750000/10775 = 10775x/10775

The fraction on the left reduces by 
dividing top and bottom by 25

    10775/431 = x 

or you may answer as a decimal 

            x = 69.60556845

(approximately).

Edwin

Answer by tutorcecilia(2152)

75=.0775x+x [Combine like-terms]
75 = 1.0775x [solve for x]
75/1.0775=1.0775x/1.0775
69.6055 = x
.
Check by plugging (x=69.6055) back into the original equation.